OpenStudy (anonymous):

Solve the rational equation (x-2)/4+x/3=3

OpenStudy (anonymous):

Could you please walk me through 2 more so I can try and become comfortable with these? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

(x-2)/4+x/3=3 12*[(x-2)/4]+12(x/3)=12*3 .... Multiply EVERY term by the LCD 12 (12(x-2))/4 + 12x/3 = 36 3(x-2) + 4x = 36 3x-6 + 4x = 36 I'll let you take over. Tell me what you get at the very end so I can check it.

OpenStudy (anonymous):

To simplify in the correct form I got 7x-42=0, is that correct?

jimthompson5910 (jim_thompson5910):

you're on the right track, what's next?

OpenStudy (anonymous):

Then I plug it into the quadratic formula

jimthompson5910 (jim_thompson5910):

no, you simply solve for x like this 7x-42=0 7x=42 x = 42/7 x = 6

jimthompson5910 (jim_thompson5910):

the quadratic formula is only used when you have a x^2 term in the equation when you're done simplifying.

OpenStudy (anonymous):

Ohh okay

OpenStudy (anonymous):

Thank you! I was was over complicating this one =)

jimthompson5910 (jim_thompson5910):

no problem, happens to me all the time :)